Rylee88:

List all perfect squares for roots +- 1 to +-15 +-1= +-15=

Olive2006:

Do you have a picture of your problem?

Rylee88:

@olive2006 wrote:
Do you have a picture of your problem?
Why would I? That's the whole problem, then it's blank.

Sailor:

Just adding this before the list, but remember that perfect square roots are always whole numbers. :^ (IM SORRY IF THIS CAUSES LAG 😭) \[\sqrt{1}=1\]\[\sqrt{4}=2\]\[\sqrt{9}=3\]\[\sqrt{16}=4\]\[\sqrt{25}=5\]\[\sqrt{36}=6\]\[\sqrt{49}=7\]\[\sqrt{64}=8\]\[\sqrt{81}=9\]\[\sqrt{100}=10\]\[\sqrt{121}=11\]\[\sqrt{144}=12\]\[\sqrt{169}=13\]\[\sqrt{196}=14\]\[\sqrt{225}=15\]

Sailor:

As you can see, if you multiply the numbers that aren't under the radical by itself, you'll get the number under the radical

Rylee88:

Positive and Negitive..

Rylee88:

Ahh, so google WAS right.

Sailor:

I forgot the negative 🗿 My apologies

Rylee88:

Yeah.